What simple interest rate will Susan need to secure to make $2,500 in interest on a $10,000 principal over 5 years?
Answer by stanbon(48558)

What simple interest rate will Susan need to secure to make $2,500 in interest on a $10,000 principal over 5 years?
-------------------------------------
I = Prt
2500 = 10000*r*5
r = 2500/50000 = 0.05 = 5%
=============================
Cheers,
Stan H.